OpenStudy (anonymous):

interval convergence of Sum of [n=0,infinity] of (1/(-9^n)*(x^(2n+1))

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ 1 }{ (-9)^{n} }\times x ^{2n+1}\]